Nameless:

On a boat, a cabin's window is in the shape of an isosceles trapezoid, as shown below. What is the area of the window? An isosceles trapezoid with top side 10 inches long and a longer bottom side. The height is 14 inches. The base of the triangle formed from the height on the right is labeled 4 inches.

Nameless:

Nameless:

pls solve

xxDeppressionxx:

a= 10 inches and b= 3+3+10=16 inches and h=15 inches. 1/2 (10+16)(15)=1/2 (26)(15) =195 sq inches

Nameless:

ty

xxDeppressionxx:

Np

jhonyy9:

@xxDeppressionxx how you get this 16 for bottom side ? pls check it

jhonyy9:

and the height is 14 in not 15 in @supie pls how is this solved ?

jhonyy9:

@darkknight pls

jhonyy9:

@supie do you see what's the problem ? bc. this is an isoscele trapezoid so the bottom side will be 10+4+4 = 18 and the height on the posted image is 14 too

supie:

It's asking for the area. I thought in order to find the area of a trapezoid I have to do the add the parallel sides together then multiply that by the height so in that case it would be \((10+8)*14\) Right?

jhonyy9:

why 8 ?

jhonyy9:

@Vocaloid pls

jhonyy9:

the bottom side not is 8

supie:

I meant 18. woops

jhonyy9:

and dont divid by 2 ?

supie:

Oh yeah, then รท2

jhonyy9:

ok but than you check these above wrote calcules these not are correct solved the posted problem

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx a= 10 inches and b= 3+3+10=16 inches and h=15 inches. 1/2 (10+16)(15)=1/2 (26)(15) =195 sq inches \(\color{#0cbb34}{\text{End of Quote}}\) look pls. here

jhonyy9:

10+16 not 16 this is 18 and the height is 14 not 15

jhonyy9:

@Nameless was helped in bad way

jhonyy9:

and now is offline

Nameless:

i online

jhonyy9:

i ve taged @xxDeppressionxx but dont wan come back and correct the posted work

Nameless:

oh ok

jhonyy9:

@Nameless ok hi your posted problem was solved not correct

jhonyy9:

please check these my above wrote details

Nameless:

ok

jhonyy9:

and than you need help pls ask me

Nameless:

ok cool

jhonyy9:

do you see what is wrong ?

Nameless:

I.T.S.

jhonyy9:

the bottom side is 18 not 16 and the height is 14 not 15

Nameless:

OH ok

Nameless:

ty

jhonyy9:

np yw

jhonyy9:

ty @supie

supie:

np